# Creation of an Fe3P Schreibersite Density Functional Tight Binding Model for Astrobiological Simulations

### v1

@@ 3.4 Molecular Dynamics
Figure 10: Radial distribution functions obtained for Fe-Fe and Fe-P interatomic distances in schreibersite, as a function of temperature and pressure. The top row (a) shows results for $T=300$K and $P=0$ GPa, while the bottom row (b) corresponds to $T=1500$K and $P=10,30,50$ GPa.
Our results for the Fe-Fe and Fe-P RDFs indicate overall very good agreement between the DFTB/ChIMES and DFT datasets (Fig. 10), where we have excluded the P-P RDFs for the sake of brevity. Simulations under ambient conditions (Fig. 10(a)) indicate that DFTB/ChIMES can effectively capture the first coordination shells, where we observe a close match with DFT for the first peak in both the Fe-Fe and Fe-P RDFs. DFTB/ChIMES reveals a slightly less structured atomic distribution for the second and third coordination shells, though the differences are relatively small. At elevated conditions (10(b)), we observe strong agreement with DFT for all solvation shells captured within our simulations. Notably, configurations at $P=50$ GPa were absent from our dataset, indicative of the ability of our model to represent state points beyond our training regime accurately.
@@ 3.5 Transferability of the model
Finally, we examine the transferability of our DFTB/ChIMES model through lattice optimizations of materials other than Fe3P schreibersite. All optimizations were performed using the material unit cells and the DFTB$+$ convergence criteria discussed above. We first performed calculations on two other iron phosphide structures, FeP and Fe2P. Iron monophosphide (FeP) belongs to a family of transition-metal monopnictides, crystallizing in the MnP structure type (orthorhombic, space group Pnma). Recently, FeP has garnered attention due to its distinctive magnetic and superconducting properties, including the formation of a helical magnetic structure74. Di-iron phosphide (Fe2), has been investigated as an anode material in sodium-ion batteries.75 Fe2P crystallizes in a hexagonal crystal structure that can be described as layers of Fe atoms in a hexagonal close-packed arrangement with P atoms occupying half of the octahedral interstices.
@@
We have also performed transferability studies on three allotropic forms of iron metal, including $\alpha$-Fe, $\gamma$-Fe, and $\epsilon$-Fe. Ferrite, or $\alpha$-iron, is the ground state and corresponds to a ferromagnetic body-centered cubic (bcc) structure that remains stable up to 1185 K. Beyond this temperature but at low pressure, $\alpha$-Fe transitions to the face-centered cubic (fcc) $\gamma$-Fe structure, also named austenite. Under pressures exceeding approximately 10-13 GPa and temperatures up to around 700 K, $\alpha$-Fe adopts the hexagonal close-packed (hcp) $\epsilon$-iron structure, also called hexaferrum77. The results of our calculations indicate strong agreement for the lattice constants of all three phases, with agreement of $\lesssim 1\%$ for both $\alpha$-Fe and $\gamma$-Fe. The agreement for $\epsilon$-Fe is also reasonable, where DFTB/ChIMES differs by $\sim$1or the $a$ lattice constant, and by $\sim$4for the $c$ lattice constant (Table 4). However, we observe some disagreement with DFT for the energetic ordering of the allotropes. In this case, DFT yields an energetic ordering of $\alpha$-Fe followed by $\epsilon$-Fe and $\gamma$-Fe, whereas DFTB/ChIMES yields $\epsilon$-Fe as the most stable phase. However, $\epsilon$-Fe is nearly isoenergetic with $\alpha$-Fe for both sets of calculations, and the energetic difference between $\alpha$-Fe and $\gamma$-Fe is similar, with a value of 0.13 eV from DFTB/ChIMES and 0.09 eV from DFT. Regardless, our results indicate the relatively strong transferability of DFTB/ChIMES to Fe/P-containing materials, given that none of these materials were present in our training set.

### v2

@@ 3.5 Molecular Dynamics
Figure 11: Radial distribution functions obtained for Fe-Fe and Fe-P interatomic distances in schreibersite, as a function of temperature and pressure. The top row (a) shows results for $T=300$K and $P=0$ GPa, while the bottom row (b) corresponds to $T=1500$K and $P=50$ GPa.
Our results for the Fe-Fe and Fe-P RDFs indicate overall very good agreement between the DFTB/ChIMES and DFT datasets (Fig. 11), where we have excluded the P-P RDFs for the sake of brevity. Simulations under ambient conditions (Fig. 11(a)) indicate that DFTB/ChIMES can effectively capture the first coordination shells, where we observe a close match with DFT for the first peak in both the Fe-Fe and Fe-P RDFs. DFTB/ChIMES reveals a slightly less structured atomic distribution for the second and third coordination shells, though the differences are relatively small. At elevated temperatures and pressures (Fig. 11(b)), we observe strong agreement with DFT for all solvation shells captured within our simulations. Notably, configurations at $P=50$ GPa were absent from our dataset, indicative of the ability of our model to accurately model state points beyond our training regime.
@@ 3.6 Transferability of the model
Finally, we examine the transferability of our DFTB/ChIMES model through lattice optimizations of materials other than Fe3P schreibersite. All optimizations were performed using the material unit cells and the DFTB$+$ convergence criteria discussed above. We first performed calculations on two other iron phosphide structures, FeP and Fe2P. Iron monophosphide (FeP) belongs to a family of transition-metal monopnictides, crystallizing in the MnP structure type (orthorhombic, space group Pnma). Recently, FeP has garnered attention due to its distinctive magnetic and superconducting properties, including the formation of a helical magnetic structure77. Di-iron phosphide (Fe2), has been investigated as an anode material in sodium-ion batteries.78 Fe2P crystallizes in a hexagonal crystal structure that can be described as layers of Fe atoms in a hexagonal close-packed arrangement with P atoms occupying half of the octahedral interstices.
@@
We have also performed transferability studies on three allotropic forms of iron metal, including $\alpha$-Fe, $\gamma$-Fe, and $\epsilon$-Fe. Ferrite, or $\alpha$-iron, is the ground state and corresponds to a ferromagnetic body-centered cubic (bcc) structure that remains stable up to 1185 K. Beyond this temperature but at low pressure, $\alpha$-Fe transitions to the face-centered cubic (fcc) $\gamma$-Fe structure, also named austenite. Under pressures exceeding approximately 10-13 GPa and temperatures up to around 700 K, $\alpha$-Fe adopts the hexagonal close-packed (hcp) $\epsilon$-iron structure, also called hexaferrum80. The results of our calculations indicate strong agreement for the lattice constants of all three phases, with agreement of $\lesssim 1\%$ for both $\alpha$-Fe and $\gamma$-Fe. The agreement for $\epsilon$-Fe is also reasonable, where DFTB/ChIMES differs by $\sim$1or the $a$ lattice constant, and by $\sim$4for the $c$ lattice constant (Table 4). However, we observe some disagreement with DFT for the energetic ordering of the allotropes. In this case, DFT yields an energetic ordering of $\alpha$-Fe followed by $\epsilon$-Fe and $\gamma$-Fe, whereas DFTB/ChIMES yields $\epsilon$-Fe as the most stable phase. However, $\epsilon$-Fe is nearly isoenergetic with $\alpha$-Fe for both sets of calculations, and the energetic difference between $\alpha$-Fe and $\gamma$-Fe is similar, with a value of 0.13 eV from DFTB/ChIMES and 0.09 eV from DFT. Regardless, our results indicate the relatively strong transferability of DFTB/ChIMES to Fe/P-containing materials, given that none of these materials were present in our training set.
@@ 4 Conclusions
Additional validation data for all sets of confining radii and ChIMES repulsive energies included in our study.
5 Data Availability
The SKF interaction files and the ChIMES parameter files for our optimal DFTB/ChIMES model are available in the following public repository: https://github.com/rdettori/dftb_FeP_Erep (last accessed November 6, 2024).